OpenStudy (anonymous):

How do you factor out x-x^5

OpenStudy (aravindg):

take x term out from both

OpenStudy (aravindg):

x(1-x^4)

OpenStudy (anonymous):

then you can also factor \(\Large (1-x^4)=(1+x^2)(1-x^2)= \) which can also be further factored with the 1-x^2...

OpenStudy (anonymous):

so do you have the COMPLETE factorization?

OpenStudy (anonymous):

x(x-1)(x+1)(x-1)(x+1)

OpenStudy (anonymous):

THANKS that helped!!!!